# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ E.3 Scaling to large datasets
In the tree construction [Lin et al., 2025], one key step involves building diffusion operators [Coifman and Lafon, 2006] between both samples and features. Although such construction can be computationally intensive, recent developments in diffusion geometry literature have introduced various techniques to significantly reduce the runtime and space complexity of this process. In particular, one can use the diffusion landmark approach [Shen and Wu, 2022], which enhances scalability by reducing the complexity from $O(n^{3})$ to $O(n^{1+2c})$ for samples and from $O(m^{3})$ to $O(m^{1+2c})$ for features, where $c<1$ represents the proportional size of the landmark set relative to the original dataset. We briefly describe the diffusion landmark approach below for the setup of the sample diffusion operator and note that it can be seamlessly applied to the feature diffusing operator. For more details on landmark diffusion, refer to [Shen and Wu, 2022].
@@ E.4 Experimental setup
Sparse approximation setup. We assess the learned hierarchical representations using a sparse approximation criterion. Specifically, we compute the $L_{1}$ norm of the Haar coefficients across all samples and features. Let $\widetilde{\mathbf{B}}_{c}$ and $\widetilde{\mathbf{B}}_{r}$ denote the Haar bases induced by the feature tree and sample tree after convergence, respectively. We measure sparsity by computing:
@@
These quantities reflect the average sparsity of the transformed data in the respective Haar bases. A lower $L_{1}$ norm indicates a more compact representation, where the signal is concentrated on fewer significant coefficients. This suggests that the learned tree structures more meaningfully represent the hierarchical organization of the data [Gavish et al., 2010, Gavish and Coifman, 2012].
Document and single-cell classification setup. For document and single-cell classification tasks (i.e., sample classification), we use the sample TWD matrix to perform classification. The dataset is split by partitioning the TWD matrix into 70% training samples and 30% testing samples. Following [Kusner et al., 2015, Huang et al., 2016], we apply a $k$-nearest neighbors classifier with $k\in\{1,3,\ldots,19\}$. The random split is repeated five times, and we report the best average classification accuracy across these runs.
Link prediction and node classification setup. We follow experimental setups used in [Chami et al., 2019, Dai et al., 2021] to maintain consistency. For the link prediction task, the set of edges is split into 85% for training, 5% for validation, and 10% for testing. A Fermi-Dirac decoder [Krioukov et al., 2010, Nickel and Kiela, 2017] is used to compute probability scores for edges, and the model is trained using cross-entropy loss with negative sampling. The link prediction performance is assessed by the area under the ROC curve (AUC). For the node classification task, dataset-specific splits are applied [Chami et al., 2019, Dai et al., 2021]: 70%/15%/15% for AIRPORT, 30%/10%/60% for DISEASE, and the standard setup of 20 training examples per class for CORA and PUBMED. The node classification is performed using a centroid-based method [Liu et al., 2019], where each class is associated with a prototype and predictions are made using a softmax classifier trained with cross-entropy loss. Additionally, an LP-based regularization objective is incorporated during training, as in [Chami et al., 2019, Dai et al., 2021]. The NC task is evaluated using the F1 score for binary-class datasets and accuracy for multi-class datasets: F1 score for DISEASE, and accuracy for AIRPORT, CORA, and PUBMED.
@@ F.1 Toy example: synthetic video recommendation system
Co-Quadtree 9.2 / 9.9 90.1$\pm$0.8 / 88.4$\pm$0.4
Co-Flowtree 9.1 / 9.9 89.2$\pm$0.7 / 89.6$\pm$0.5
Co-WCTWD 9.3 / 9.8 82.6$\pm$1.0 / 86.4$\pm$0.8
Co-WQTWD 9.3 / 9.7 83.1$\pm$0.8 / 85.9$\pm$0.7
Co-UltraTree 8.9 / 9.0 88.5$\pm$1.4 / 82.3$\pm$0.9
Co-TSWD-1 10.3 / 11.4 85.6$\pm$1.4 / 83.2$\pm$1.1
Co-TSWD-5 10.1 / 10.5 87.2$\pm$0.8 / 84.9$\pm$0.9
Co-TSWD-10 9.7 / 10.0 89.7$\pm$0.8 / 87.4$\pm$0.8
Co-SWCTWD 8.2 / 8.9 89.4$\pm$1.0 / 88.3$\pm$0.4
Co-SWQTWD 8.4 / 8.6 88.6$\pm$0.2 / 88.4$\pm$0.9
Co-MST-TWD 10.5 / 10.9 82.1$\pm$1.7 / 80.6$\pm$1.3
Co-TR-TWD 9.8 / 10.2 92.6$\pm$0.4 / 93.8$\pm$0.7
Co-HHC-TWD 10.3 / 11.4 89.2$\pm$2.0 / 88.1$\pm$1.7
Co-gHHC-TWD 10.1 / 10.9 88.4$\pm$1.6 / 87.0$\pm$1.3
Co-UltraFit-TWD 9.7 / 10.3 90.1$\pm$1.3 / 89.5$\pm$1.2
QUE 7.6 / 8.1 93.4$\pm$1.7 / 92.5$\pm$0.5
WSV - 92.5$\pm$0.8 / 91.6$\pm$0.6
Tree-WSV 8.4 / 8.7 93.9$\pm$0.6 / 92.7$\pm$0.8
Alg. 1 3.2 / 4.0 96.2$\pm$0.3 / 95.3$\pm$0.4
Alg. 2 2.5 / 3.1 99.4$\pm$0.2 / 99.6$\pm$0.1
Note that other functions of the distance can also be considered, e.g., elastic potential operator [Coifman and Gavish, 2011]. In our synthetic experiment, we set $\sigma_{c}=1$, $\sigma_{r}=0.6$, $\sigma_{0}=0.5$, $\sigma_{1}=0.25$, $\sigma_{\text{noise}}$, and $d=30$. Finally, to simulate an unstructured observation, we apply random permutations [Mishne et al., 2017, Yair et al., 2017] to the rows and columns of $\mathbf{Y}$ to obtain the observed data matrix $\mathbf{X}$.
@@ F.2 Empirical uniqueness of hierarchical representations under strong regularization
For sufficiently large $\gamma_{r}$ and $\gamma_{c}$, the resulting TWDs are unique. This means that even if the initial distances are random, our methods will converge to unique trees and TWDs. Fig. 7 illustrates that using different random initializations yields identical final expansion coefficients for the Zeisel dataset. However, we would like to point out that such large regularization does not necessarily lead to the best hierarchical representation. When $\gamma_{r}$ and $\gamma_{c}$ are too large, the regularization terms dominate the distance computation in the iterative learning scheme, leading to trees that may lack meaningful hierarchical structure. This is evident when comparing these results to those in Fig. 2 in Sec. 6, where moderate regularization (tuned by Optuna [Akiba et al., 2019]) yields more informative trees. The purpose of this experiment is therefore to demonstrate empirical uniqueness under strong regularization, rather than to advocate for such settings in practice.
@@ F.3 Ablation study: effectiveness of iterative learning scheme in classification tasks
Table 6: Document and single-cell classification accuracy using the initial TWD (without iterative refinement) and the updated TWD (with iterative refinement). Results marked with * are taken from [Lin et al., 2025]. Arrows ($\uparrow$) indicate improvements achieved through the iterative learning process. The best performance is marked in bold, and the second-best is underlined.
@@
Quadtree 95.5$\pm$0.5 69.6$\pm$0.8 95.9$\pm$0.4 89.3$\pm$0.3 80.1$\pm$1.2 80.6$\pm$0.6
Co-Quadtree 96.2$\pm$0.4 ($\uparrow$) 69.6$\pm$0.3 95.9$\pm$0.2 89.4$\pm$0.2 ($\uparrow$) 81.7$\pm$1.0 ($\uparrow$) 80.7$\pm$0.3 ($\uparrow$)
Flowtree 95.3$\pm$1.1 70.2$\pm$0.9 94.4$\pm$0.6 90.1$\pm$0.3 81.7$\pm$0.9 81.8$\pm$0.9
Co-Flowtree 95.7$\pm$0.9 ($\uparrow$) 71.5$\pm$0.7 ($\uparrow$) 95.6$\pm$0.5 ($\uparrow$) 91.4$\pm$0.4 ($\uparrow$) 84.3$\pm$0.7 ($\uparrow$) 83.0$\pm$1.2 ($\uparrow$)
WCTWD 92.6$\pm$2.1 69.1$\pm$2.6 93.7$\pm$2.9 88.2$\pm$1.4 81.3$\pm$4.9 78.4$\pm$3.3
Co-WCTWD 93.2$\pm$1.2 ($\uparrow$) 70.2$\pm$2.1 ($\uparrow$) 94.7$\pm$2.6 ($\uparrow$) 87.4$\pm$1.0 82.5$\pm$2.9 ($\uparrow$) 79.4$\pm$2.1 ($\uparrow$)
WQTWD 94.3$\pm$1.7 69.4$\pm$2.4 94.6$\pm$3.2 87.4$\pm$1.8 80.9$\pm$3.5 79.1$\pm$3.0
Co-WQTWD 95.7$\pm$1.8 ($\uparrow$) 70.7$\pm$2.2 ($\uparrow$) 95.5$\pm$1.3 ($\uparrow$) 88.2$\pm$2.1 ($\uparrow$) 82.3$\pm$3.1 ($\uparrow$) 80.5$\pm$2.8 ($\uparrow$)
UltraTree 93.1$\pm$1.5 68.1$\pm$3.2 92.3$\pm$1.9 86.2$\pm$3.1 83.9$\pm$1.6 82.3$\pm$2.6
Co-UltraTree 95.3$\pm$1.4 ($\uparrow$) 70.1$\pm$2.8 ($\uparrow$) 93.6$\pm$2.0 ($\uparrow$) 86.5$\pm$2.8 ($\uparrow$) 85.8$\pm$1.1 ($\uparrow$) 84.6$\pm$1.3 ($\uparrow$)
TSWD-1 87.6$\pm$1.9 69.8$\pm$1.3 94.5$\pm$0.5 85.5$\pm$0.6 79.6$\pm$1.8 72.6$\pm$1.8
Co-TSWD-1 88.2$\pm$1.4 ($\uparrow$) 70.4$\pm$1.2 ($\uparrow$) 94.7$\pm$0.9 ($\uparrow$) 86.1$\pm$0.5 ($\uparrow$) 80.2$\pm$1.4 ($\uparrow$) 73.2$\pm$1.0
TSWD-5 88.1$\pm$1.3 70.5$\pm$1.1 95.9$\pm$0.4 90.8$\pm$0.1 81.3$\pm$1.4 74.9$\pm$1.1
Co-TSWD-5 88.7$\pm$1.7 ($\uparrow$) 71.0$\pm$1.5 ($\uparrow$) 96.7$\pm$0.8 ($\uparrow$) 91.5$\pm$0.4 ($\uparrow$) 82.0$\pm$0.9 ($\uparrow$) 75.4$\pm$0.7 ($\uparrow$)
TSWD-10 88.6$\pm$0.9 70.7$\pm$1.3 95.9$\pm$0.6 91.1$\pm$0.5 83.2$\pm$0.8 76.5$\pm$0.7
Co-TSWD-10 89.2$\pm$1.1 ($\uparrow$) 71.4$\pm$1.8 ($\uparrow$) 95.5$\pm$0.2 91.8$\pm$0.7 ($\uparrow$) 83.8$\pm$0.5 ($\uparrow$) 77.2$\pm$0.9 ($\uparrow$)
SWCTWD 92.8$\pm$1.2 70.2$\pm$1.2 94.1$\pm$1.8 90.2$\pm$1.2 81.9$\pm$3.1 78.3$\pm$1.7
Co-SWCTWD 93.5$\pm$2.4 ($\uparrow$) 70.5$\pm$1.0 ($\uparrow$) 94.4$\pm$1.3 ($\uparrow$) 90.7$\pm$1.5 ($\uparrow$) 82.7$\pm$1.7 ($\uparrow$) 79.0$\pm$0.9 ($\uparrow$)
SWQTWD 94.5$\pm$1.0 70.6$\pm$1.9 95.4$\pm$2.0 89.8$\pm$1.1 80.7$\pm$2.5 79.8$\pm$2.5
Co-SWQTWD 96.2$\pm$1.2 ($\uparrow$) 72.4$\pm$2.1 ($\uparrow$) 96.0$\pm$1.1 ($\uparrow$) 90.6$\pm$2.3 ($\uparrow$) 82.4$\pm$1.4 ($\uparrow$) 81.3$\pm$1.1 ($\uparrow$)
MST-TWD 88.4$\pm$1.9 68.2$\pm$1.9 90.0$\pm$3.1 86.4$\pm$1.2 80.1$\pm$3.1 76.2$\pm$2.5
Co-MST-TWD 88.7$\pm$2.4 ($\uparrow$) 68.4$\pm$3.3 ($\uparrow$) 91.3$\pm$2.9 ($\uparrow$) 87.1$\pm$1.4 ($\uparrow$) 80.1$\pm$2.8 76.5$\pm$1.3 ($\uparrow$)
TR-TWD 89.2$\pm$0.9 70.2$\pm$0.7 92.9$\pm$0.8 88.7$\pm$1.1 80.3$\pm$0.7 78.4$\pm$1.2
Co-TR-TWD 89.5$\pm$1.2 ($\uparrow$) 70.9$\pm$1.7 ($\uparrow$) 93.4$\pm$2.2 ($\uparrow$) 89.5$\pm$1.4 ($\uparrow$) 80.7$\pm$0.8 ($\uparrow$) 78.5$\pm$0.9 ($\uparrow$)
HHC-TWD 85.3$\pm$1.8 70.4$\pm$0.4 93.4$\pm$0.8 88.5$\pm$0.7 82.3$\pm$0.7 77.3$\pm$1.1
Co-HHC-TWD 86.1$\pm$2.1 ($\uparrow$) 70.1$\pm$1.3 ($\uparrow$) 93.6$\pm$1.5 ($\uparrow$) 88.5$\pm$0.5 83.2$\pm$1.4 ($\uparrow$) 77.6$\pm$0.8 ($\uparrow$)
gHHC-TWD 83.2$\pm$2.4 69.9$\pm$1.8 90.3$\pm$2.2 86.9$\pm$2.0 79.4$\pm$1.9 73.6$\pm$1.6
Co-gHHC-TWD 84.0$\pm$2.0 ($\uparrow$) 70.4$\pm$1.6 ($\uparrow$) 90.7$\pm$1.7 ($\uparrow$) 87.2$\pm$1.9 ($\uparrow$) 79.9$\pm$1.4 ($\uparrow$) 84.2$\pm$1.2 ($\uparrow$)
UltraFit-TWD 84.9$\pm$1.4 69.5$\pm$1.2 91.6$\pm$0.9 87.4$\pm$1.6 81.9$\pm$3.3 77.8$\pm$1.2
Co-UltraFit-TWD 86.8$\pm$0.9 ($\uparrow$) 70.9$\pm$1.1 ($\uparrow$) 91.9$\pm$1.0 ($\uparrow$) 89.9$\pm$2.0 ($\uparrow$) 83.7$\pm$2.9 ($\uparrow$) 79.1$\pm$1.8 ($\uparrow$)
WMD 95.4$\pm$0.7 71.3$\pm$0.6 97.2$\pm$0.1 92.6$\pm$0.3 - -
GMD - - - - 84.2$\pm$0.7 81.4$\pm$0.7
HD-TWD 96.1$\pm$0.4 73.4$\pm$0.2 96.9$\pm$0.2 93.1$\pm$0.4 89.1$\pm$0.4 84.3$\pm$0.3
Alg. 1 96.7$\pm$0.3 ($\uparrow$) 74.1$\pm$0.5 ($\uparrow$) 97.3$\pm$0.2 ($\uparrow$) 94.0$\pm$0.4 90.1$\pm$0.4 ($\uparrow$) 86.7$\pm$0.5 ($\uparrow$)
Alg. 2 97.3$\pm$0.5 ($\uparrow$) 76.7$\pm$0.7 ($\uparrow$) 97.6$\pm$0.1 ($\uparrow$) 94.2$\pm$0.2 ($\uparrow$) 94.0$\pm$0.6 ($\uparrow$) 93.3$\pm$0.7 ($\uparrow$)
@@ F.4 Ablation study: incorporating Wavelet filtering with alternative TWD methods
The key difference between Alg. 1 and Alg. 2 is the application of a filtering step in each iteration of Alg. 2, which aims to suppress noise and other nuisance components. This raises the question of whether wavelet filtering could
Table 7: The $L_{1}$ norm of the Haar coefficients across all samples and features when applying wavelet filtering induced by trees constructed from various TWD-based baselines. Lower values indicate sparser representations. Arrows ()($\uparrow$)( ↑ ) denote cases where wavelet filtering leads to an improvement (i.e., reduction in $L_{1}$ norm) compared to the unfiltered baseline. Values are reported in the format $\mathrm{samples}$ / $\mathrm{features}$ (the lowest in bold and the second lowest underlined).
@@
also benefit other TWD-based baseline methods when considering in our iterative learning scheme. To explore this, we apply Haar wavelet filtering using trees constructed from various TWD baselines and report the resulting $L_{1}$ norms of the Haar coefficients across all samples and features in Tab. 7. For each baseline, we append “-Wavelet” to indicate the variant with wavelet filtering applied. We observe that incorporating wavelet filtering into these baselines does not consistently lead to improved hierarchical representations. We argue that this is because the trees used in these baselines are primarily optimized to approximate Wasserstein distances as ground metrics, rather than to represent the hierarchical structure of the data. As such, they do not serve as meaningful hierarchical representations, since the Wasserstein metric is inherently not a hierarchical metric. Consequently, applying wavelet filters, whose effectiveness depends on the data and the tree structure, does not improve, and may even hinder, hierarchical representation learning.
We further evaluate the effect of wavelet filtering on downstream classification tasks for document and single-cell datasets. Tab. 8 reports the classification accuracies obtained by applying wavelet filtering within the iterative learning scheme using alternative TWD methods. Similarly, for each baseline, we append “-Wavelet” to indicate the variant
Table 8: Classification accuracy on document and single-cell datasets using various TWD-based baselines with and without wavelet filtering, applied within the iterative learning scheme. The “-Wavelet” suffix denotes the inclusion of Haar wavelet filtering. Arrows ()($\uparrow$)( ↑ ) denote cases where wavelet filtering leads to an improvement compared to the unfiltered baseline. While some baselines show marginal improvement, the gains are inconsistent across datasets, highlighting the importance of using meaningful hierarchical structures for wavelet filtering to enhance representation learning. The best performance is marked in bold, and the second-best is underlined.
@@
Co-Quadtree 96.2$\pm$0.4 69.6$\pm$0.3 95.9$\pm$0.2 89.4$\pm$0.2 81.7$\pm$1.0 80.7$\pm$0.3
Co-Quadtree-Wavelet 95.6$\pm$0.6 69.6$\pm$0.2 96.0$\pm$0.1 ($\uparrow$) 89.4$\pm$1.0 81.4$\pm$1.1 80.6$\pm$0.4 ($\uparrow$)
Co-Flowtree 95.7$\pm$0.9 71.5$\pm$0.7 95.6$\pm$0.5 91.4$\pm$0.4 84.3$\pm$0.7 83.0$\pm$1.2
Co-Flowtree-Wavelet 95.5$\pm$0.8 71.0$\pm$ 0.8 95.0$\pm$0.7 90.8$\pm$0.5 82.7$\pm$0.8 82.2$\pm$1.0
Co-WCTWD 93.2$\pm$1.2 70.2$\pm$2.1 94.7$\pm$2.6 87.4$\pm$1.0 82.5$\pm$2.9 79.4$\pm$2.1
Co-WCTWD-Wavelet 93.2$\pm$1.1 70.6$\pm$1.9 ($\uparrow$) 94.5$\pm$2.2 88.3$\pm$1.3 ($\uparrow$) 82.2$\pm$3.0 79.0$\pm$2.9
Co-WQTWD 95.7$\pm$1.8 70.7$\pm$2.2 95.5$\pm$1.3 88.2$\pm$2.1 82.3$\pm$3.1 80.5$\pm$2.8
Co-WQTWD-Wavelet 95.0$\pm$1.6 70.2$\pm$1.8 95.2$\pm$ 1.8 88.2$\pm$1.9 81.9$\pm$3.0 80.7$\pm$2.7 ($\uparrow$)
Co-UltraTree 95.3$\pm$1.4 70.1$\pm$2.8 93.6$\pm$2.0 86.5$\pm$2.8 85.8$\pm$1.1 84.6$\pm$1.3
Co-UltraTree-Wavelet 95.0$\pm$1.5 70.4$\pm$2.9 ($\uparrow$) 93.0$\pm$1.8 86.3$\pm$2.7 85.7$\pm$1.3 83.9$\pm$2.0
Co-TSWD-1 88.2$\pm$1.4 70.4$\pm$1.2 94.7$\pm$0.9 86.1$\pm$0.5 80.2$\pm$1.4 73.2$\pm$1.0
Co-TSWD-1-Wavelet 88.0$\pm$1.3 70.1$\pm$1.2 94.6$\pm$1.0 85.7$\pm$0.7 80.1$\pm$1.2 72.9$\pm$1.1
Co-TSWD-5 88.7$\pm$1.7 71.0$\pm$1.5 96.7$\pm$0.8 91.5$\pm$0.4 82.0$\pm$0.9 75.4$\pm$0.7
Co-TSWD-5-Wavelet 88.5$\pm$1.4 70.9$\pm$1.3 96.5$\pm$0.7 91.1$\pm$0.6 82.3$\pm$1.0 ($\uparrow$) 75.6$\pm$0.9 ($\uparrow$)
Co-TSWD-10 89.2$\pm$1.1 71.4$\pm$1.8 95.5$\pm$0.2 91.8$\pm$0.7 83.8$\pm$0.5 77.2$\pm$0.9
Co-TSWD-10-Wavelet 89.0$\pm$0.9 71.2$\pm$1.4 95.4$\pm$0.5 91.3$\pm$0.8 83.7$\pm$0.7 76.9$\pm$1.0
Co-SWCTWD 93.5$\pm$2.4 70.5$\pm$1.0 94.4$\pm$1.3 90.7$\pm$1.5 82.7$\pm$1.7 79.0$\pm$0.9
Co-SWCTWD-Wavelet 93.2$\pm$1.6 71.3$\pm$0.9 ($\uparrow$) 94.2$\pm$1.5 90.3$\pm$1.7 82.3$\pm$2.0 79.0$\pm$0.5
Co-SWQTWD 96.2$\pm$1.2 72.4$\pm$2.1 96.0$\pm$1.1 90.6$\pm$2.3 82.4$\pm$1.4 81.3$\pm$1.1
Co-SWQTWD-Wavelet 95.9$\pm$0.9 71.8$\pm$1.6 95.6$\pm$1.3 90.2$\pm$2.0 82.0$\pm$1.7 80.9$\pm$1.4
Co-MST-TWD 88.7$\pm$2.4 68.4$\pm$3.3 91.3$\pm$2.9 87.1$\pm$1.4 80.1$\pm$2.8 76.5$\pm$1.3
Co-MST-TWD-Wavelet 88.6$\pm$1.3 68.7$\pm$2.9 ($\uparrow$) 90.8$\pm$2.7 86.9$\pm$1.3 80.2$\pm$2.6 ($\uparrow$) 76.1$\pm$2.1
Co-TR-TWD 89.5$\pm$1.2 70.9$\pm$1.7 93.4$\pm$2.2 89.5$\pm$1.4 80.7$\pm$0.8 78.5$\pm$0.9
Co-TR-TWD-Wavelet 90.6$\pm$0.7 ($\uparrow$) 70.9$\pm$1.3 93.2$\pm$1.9 89.0$\pm$1.3 80.5$\pm$0.9 78.5$\pm$0.7
Co-HHC-TWD 86.1$\pm$2.1 70.1$\pm$1.3 93.6$\pm$1.5 88.5$\pm$0.5 83.2$\pm$1.4 77.6$\pm$0.8
Co-HHC-TWD-Wavelet 85.9$\pm$1.7 70.4$\pm$1.2 ($\uparrow$) 93.4$\pm$1.1 89.2$\pm$0.9 ($\uparrow$) 82.7$\pm$1.0 77.5$\pm$0.9
Co-gHHC-TWD 84.0$\pm$2.0 70.4$\pm$1.6 90.7$\pm$1.7 87.2$\pm$1.9 79.9$\pm$1.4 84.2$\pm$1.2
Co-gHHC-TWD-Wavelet 83.5$\pm$1.7 70.2$\pm$1.9 90.6$\pm$2.1 87.0$\pm$1.4 82.7$\pm$2.5 78.6$\pm$1.6
Co-UltraFit-TWD 86.8$\pm$0.9 70.9$\pm$1.1 91.9$\pm$1.0 89.9$\pm$2.0 83.7$\pm$2.9 79.1$\pm$1.8
Co-UltraFit-TWD-Wavelet 86.7$\pm$1.2 70.3$\pm$1.4 92.1$\pm$0.9 ($\uparrow$) 88.7$\pm$1.8 83.4$\pm$2.7 79.0$\pm$1.7
QUE 84.7$\pm$0.5 72.4$\pm$0.6 91.9$\pm$0.5 91.6$\pm$0.9 83.6$\pm$1.4 82.5$\pm$1.9
QUE-Wavelet 86.4$\pm$0.7 ($\uparrow$) 72.0$\pm$0.8 92.3$\pm$0.8 ($\uparrow$) 90.8$\pm$1.0 85.1$\pm$1.0 ($\uparrow$) 82.4$\pm$1.6
WSV 85.9$\pm$1.0 71.4$\pm$1.3 92.6$\pm$0.7 89.0$\pm$1.5 81.6$\pm$2.4 77.5$\pm$1.7
Tree-WSV 86.3$\pm$1.5 71.2$\pm$1.9 92.4$\pm$1.0 88.7$\pm$1.9 82.0$\pm$2.9 76.4$\pm$2.4
Tree-WSV-Wavelet 86.2$\pm$1.1 71.5$\pm$1.4 ($\uparrow$) 92.0$\pm$0.8 89.3$\pm$1.4 ($\uparrow$) 82.0$\pm$2.3 78.1$\pm$1.9 ($\uparrow$)
Alg. 1 96.7$\pm$0.3 74.1$\pm$0.5 97.3$\pm$0.2 94.0$\pm$0.4 90.1$\pm$0.4 86.7$\pm$0.5
Alg. 2 97.3$\pm$0.5 ($\uparrow$) 76.7$\pm$0.7 ($\uparrow$) 97.6$\pm$0.1 ($\uparrow$) 94.2$\pm$0.2 ($\uparrow$) 94.0$\pm$0.6 ($\uparrow$) 93.3$\pm$0.7 ($\uparrow$)
with wavelet filtering applied. We see that the results show inconsistent improvements across datasets and domains, with any gains being marginal at best. In contrast, when wavelet filtering is applied within our proposed method, the improvement is significantly more pronounced. These findings are consistent with the sparse approximation results in Tab. 7, reinforcing the observation that wavelet filtering effectively improves hierarchical representation learning only when applied to trees that meaningfully represent the data hierarchy.

### v3

@@ E.3 Scaling to large datasets
In the tree construction [37], one key step involves building diffusion operators [51] between both samples and features. Although such construction can be computationally intensive, recent developments in diffusion geometry literature have introduced various techniques to significantly reduce the runtime and space complexity of this process. In particular, one can use the diffusion landmark approach [76], which enhances scalability by reducing the complexity from $O(n^{3})$ to $O(n^{1+2c})$ for samples and from $O(m^{3})$ to $O(m^{1+2c})$ for features, where $c<1$ represents the proportional size of the landmark set relative to the original dataset. We briefly describe the diffusion landmark approach below for the setup of the sample diffusion operator and note that it can be seamlessly applied to the feature diffusing operator.
@@ E.4 Experimental setup
Sparse approximation setup.
We assess the learned hierarchical representations using a sparse approximation criterion. Specifically, we compute the $L_{1}$ norm of the Haar coefficients across all samples and features. Let $\widetilde{\mathbf{B}}_{c}$ and $\widetilde{\mathbf{B}}_{r}$ denote the Haar bases induced by the feature tree and sample tree after convergence, respectively. We measure sparsity by computing:
@@
These quantities reflect the average sparsity of the transformed data in the respective Haar bases. A lower $L_{1}$ norm indicates a more compact representation, where the signal is concentrated on fewer significant coefficients. This suggests that the learned tree structures more meaningfully represent the hierarchical organization of the data [43, 53].
Document and single-cell classification setup.
For document and single-cell classification tasks (i.e., sample classification), we use the sample TWD matrix to perform classification. The dataset is split by partitioning the TWD matrix into 70% training samples and 30% testing samples. We apply a $k$-nearest neighbors classifier with $k\in\{1,3,\ldots,19\}$. The random split is repeated five times [25, 26], and we report the best average classification accuracy across these runs.
Link prediction and node classification setup.
We follow experimental setups [21, 45] to maintain consistency. For the link prediction task, the set of edges is split into 85% for training, 5% for validation, and 10% for testing. A Fermi-Dirac decoder [93, 13] is used to compute probability scores for edges, and the model is trained using cross-entropy loss with negative sampling. The link prediction performance is assessed by the area under the ROC curve (AUC). For the node classification task, dataset-specific splits are applied [21, 45]: 70%/15%/15% for AIRPORT, 30%/10%/60% for DISEASE, and the standard setup of 20 training examples per class for CORA and PUBMED. The node classification is performed using a centroid-based method [94], where each class is associated with a prototype and predictions are made using a softmax classifier trained with cross-entropy loss. Additionally, an LP-based regularization objective is incorporated during training [21, 45]. The NC task is evaluated using the F1 score for binary-class datasets and accuracy for multi-class datasets: F1 score for DISEASE, and accuracy for AIRPORT, CORA, and PUBMED.
@@ F.1 Toy example: synthetic video recommendation system
Note that other functions of the distance can also be considered, e.g., elastic potential operator [154]. In our synthetic experiment, we set $\sigma_{c}=1$, $\sigma_{r}=0.6$, $\sigma_{0}=0.5$, $\sigma_{1}=0.25$, $\sigma_{\text{noise}}$, and $d=30$. Finally, to simulate an unstructured observation, we apply random permutations [129, 54] to the rows and columns of $\mathbf{Y}$ to obtain the observed data matrix $\mathbf{X}$.
@@
[…]
@@ F.2 Empirical uniqueness of hierarchical representations under strong regularization
For sufficiently large $\gamma_{r}$ and $\gamma_{c}$, the resulting TWDs are unique. This means that even if the initial distances are random, our methods will converge to unique trees and TWDs. Fig. 7 illustrates that using different random initializations yields identical final expansion coefficients for the Zeisel dataset. However, we would like to point out that such large regularization does not necessarily lead to the best hierarchical representation. When $\gamma_{r}$ and $\gamma_{c}$ are too large, the regularization terms dominate the distance computation in the iterative learning scheme, leading to trees that may lack meaningful hierarchical structure. This is evident when comparing these results to those in Fig. 2 in Sec. 6, where moderate regularization (tuned by Optuna [151]) yields more informative trees. The purpose of this experiment is therefore to demonstrate empirical uniqueness under strong regularization, rather than to advocate for such settings in practice.
@@ F.3 Ablation study: effectiveness of iterative learning scheme in classification tasks
Table 6: Document and single-cell classification accuracy using the initial TWD (without iterative refinement) and the updated TWD (with iterative refinement). Results marked with * are taken from the work of Lin et al. [37]. Arrows ($\uparrow$) indicate improvements achieved through the iterative learning process. The best performance is marked in bold, and the second-best is underlined.
@@
[…]
@@ F.4 Ablation study: incorporating Wavelet filtering with alternative TWD methods
The key difference between Alg. 1 and Alg. 2 is the application of a filtering step in each iteration of Alg. 2, which aims to suppress noise and other nuisance components. This raises the question of whether wavelet filtering could also benefit other TWD-based baseline methods when considering in our iterative learning scheme. To explore this, we apply Haar wavelet filtering using trees constructed from various TWD baselines and report the resulting $L_{1}$ norms of the Haar coefficients across all samples and features in Tab. 7. For each baseline, we append “-Wavelet” to indicate the variant with wavelet filtering applied. We observe that incorporating wavelet filtering into these baselines does not consistently lead to improved hierarchical representations. We argue that this is because the trees used in these baselines are primarily optimized to approximate Wasserstein distances as ground metrics, rather than to represent the hierarchical structure of the data. As such, they do not serve as meaningful hierarchical representations, since the Wasserstein metric is inherently not a hierarchical metric. Consequently, applying wavelet filters, whose effectiveness depends on the data and the tree structure, does not improve, and may even hinder, hierarchical representation learning.
We further evaluate the effect of wavelet filtering on downstream classification tasks for document and single-cell datasets. Tab. 8 reports the classification accuracies obtained by applying wavelet filtering within the iterative learning scheme using alternative TWD methods. Similarly, for each baseline, we append “-Wavelet” to indicate the variant with wavelet filtering applied. We see that the results show inconsistent improvements across datasets and domains, with any gains being marginal at best. In contrast, when wavelet filtering is applied within our proposed method, the improvement is significantly more pronounced. These findings are consistent with the sparse approximation results in Tab. 7, reinforcing the observation that wavelet filtering effectively improves hierarchical representation learning only when applied to trees that meaningfully represent the data hierarchy.
Table 7: The $L_{1}$ norm of the Haar coefficients across all samples and features when applying wavelet filtering induced by trees constructed from various TWD-based baselines. Lower values indicate sparser representations. Arrows $(\uparrow)$ denote cases where wavelet filtering leads to an improvement (i.e., reduction in $L_{1}$ norm) compared to the unfiltered baseline. Values are reported in the format $\mathrm{samples}$ / $\mathrm{features}$ (the lowest in bold and the second lowest underlined).
@@
Table 8: Classification accuracy on document and single-cell datasets using various TWD-based baselines with and without wavelet filtering, applied within the iterative learning scheme. The “-Wavelet” suffix denotes the inclusion of Haar wavelet filtering. Arrows $(\uparrow)$ denote cases where wavelet filtering leads to an improvement compared to the unfiltered baseline. While some baselines show marginal improvement, the gains are inconsistent across datasets, highlighting the importance of using meaningful hierarchical structures for wavelet filtering to enhance representation learning. The best performance is marked in bold, and the second-best is underlined.
@@
[…]